# An elementary method to determine the critical mass of a sphere of fissile material based on a separation of neutron transport and nuclear reaction processes

### v1

@@ II.1 Determination of $\ell$ from nuclear reaction cross sections
    $N$ Neutrons will leave the sphere, according to the definition of $\tau_{w}$ and $\ell$.
@@
The neutron emission rate $N/\tau_{w}$ from the sphere is reflected by $-N$ in the above equation.
It might seem strange that $N$ is a constant and that the emission or “leakage rate remains constant in the presence of other processes. But this is simply as expected with $N$ specified; From kinetic theory, the current $I$ through the surface is $I=N\widebar{v}A/4V$ where $\widebar{v}$ is the average velocity, $A$ is the surface area and $V$ is the volume, and by our definitions $N=I\tau_{w}$ as $\dot{N}/N=1/\tau_{w}$. The presence of other processes does not affect this “leakage” rate. Alternatively, the entire problem could be cast in terms of rate equations; however, there is no pressing need to do that.
@@ II.2 Neutron random walk
In Appendix A, it is shown that the average squared distance from any point within a sphere of radius $R$ to the surface is $54R^{2}/$ 5. As shown in Appendix B, there is a modest correction factor because the square root of the average squared displacement from a random walk overestimates the actual average displacement. The factor of 4/5 and this correction factor can be combined together into a dimensionless scale factor $\epsilon$, so we write $\ell_{0}^{2}=R^{2}/\epsilon^{2}$ and cast Eq. (9) as
@@ II.3 Neutron travel distance to escape and the critical radius
$A$ Atomic weight; mass number
@@
$N$ Number of neutrons in the sphere
@@ III Applications
[…]
@@ V Neutron production probability
In the case of a subcritical sphere, that is, when the right hand side of Eq. (6) is $11-\eta<$ 1, the number $N$ will decay over time, but over a much longer time $\tau=\tau_{w}/\eta$. If $N_{i}$ neutrons are introduced into the sphere, the total number emitted is
[…]
@@ VI Fission Time Evolution
The number of generations $N_{g}$ to cause a mole of fissions when starting from a single neutron is then $2.13^{N_{g}}=6.02\times 10^{23}$, or $N_{g}\sim 72$. The mean distance traveled between fissions is $1/n\sigma_{f}=11\ {\rm cm}$. (Scattering occurs over a smaller distance, but this does not matter so far as computing a timescale is concerned.) The total distance traveled is consequently $72\times 11\sim 800$ cm for the uncompressed metal. The velocity of a 1 MeV neutron is $1.4\times 10^{9}$ cm/s, so the time scale is $800/1.4\times 10^{9}=0.57\times 10^{-6}$ s or 0.57 $\mu$s. Nuclear explosions are microsecond-scale events.
@@ Appendix A Average squared distance to sphere surface
[…]
The odd-parity term in $\mathrm{d}^{2}(\theta,z)$ with $\cos\theta$ does not contribute to the integral over $\int_{0}^{\pi}\sin\theta\ d\theta$. Averaging over $z$ must include weighting by the volume of a spherical shell of radius $z$ and thickness $dz$ within a sphere of radius unity and volume $34\pi/$ 3:
[…]
@@ Appendix C Estimation of the effect of forward scattering on the mean free path
[…]
@@ Appendix D Application to the diffusion equation
$-D\left[\nabla_{r}n(r)\right]\Big{|}_{r=R}=\frac{1}{4}n_{g}\widebar{v}$ (29)
@@
[…]
@@
$-\ell^{\prime}+\frac{4}{3}x+\frac{6}{\pi}x^{2}=0$ (36)

### v2

# arXiv:2506.06927v2[physics.pop-ph] 01 Mar 2026

@@ II.1 Determination of $\ell$ from nuclear reaction cross sections
    NN Neutrons will leave the sphere, according to the definition of $\tau_{w}$ and $\ell$.
@@
The neutron emission rate $N/\tau_{w}$ from the sphere is reflected by N-N in the above equation.
It might seem strange that $N$ is a constant and that the emission or “leakage rate remains constant in the presence of other processes. But this is simply as expected with $N$ specified; From kinetic theory, the current $I$ through the surface is $I=N\widebar{v}A/4V$ where $\widebar{v}$ is the average velocity, $A$ is the surface area and $V$ is the volume, and by our definitions $N=I\tau_{w}$ as $\dot{N}/N=1/\tau_{w}$. The presence of other processes does not affect this “leakage” rate. Alternatively, the entire problem could be cast in terms of rate equations (see Appendix F); however, there is no pressing need to do that.
@@ II.2 Neutron random walk
In Appendix A, it is shown that the average squared distance from any point within a sphere of radius $R$ to the surface is $4R^{2}/5$. As shown in Appendix B, there is a modest correction factor because the square root of the average squared displacement from a random walk overestimates the actual average displacement. The factor of 4/5 and this correction factor can be combined together into a dimensionless scale factor $\epsilon$, so we write $\ell_{0}^{2}=R^{2}/\epsilon^{2}$ and cast Eq. (9) as
@@ II.3 Neutron travel distance to escape and the critical radius
AA Atomic weight; mass number
@@
NN Number of neutrons in the sphere
@@ III Applications
At an enrichment of 19.75%, with $\sigma_{f}=.308$, a critical mass of 502.8 kg is significantly less than the 800 kg cited in ref. [10] for 20%. With $\sigma_{f}=0.1$, a critical mass is not possible at enrichment of 20%. However, one can find differing values in the literature for lower enrichment levels. In a 2006 paper, Glaser examined the proliferation potential of uranium enriched to various levels using the then-current MCNP 4C code, estimating a critical mass of 780 kg for 19.75%.[12] We could adjust our guess of $\sigma_{f}=0.1$ to bring the values into better agreement, however, it is more instructive to retain our original estimate. One implication regarding our estimate is that neutron moderation is not as significant as one might expect, because neutrons are constantly being absorbed and replaced by “fresh” fission-produced neutrons with a pristine energy distribution.
@@ V Neutron production probability
In the case of a subcritical sphere, that is, when the right hand side of Eq. (6) is $1-\eta<1$, the number $N$ will decay over time, but over a much longer time $\tau=\tau_{w}/\eta$. If $N_{i}$ neutrons are introduced into the sphere, the total number emitted is
$N_{e}=\frac{N_{i}}{\tau_{w}}\int_{0}^{\infty}e^{-\eta t/\tau_{w}}dt=\frac{N_{i}}{\eta}.$ (18)
@@ VI Fission Time Evolution
[…]
@@ Appendix A Average squared distance to sphere surface
$\widebar{\mathrm{d}^{2}}(z)=\frac{1}{4\pi}\int d^{2}(z,\theta)d\Omega=1-\frac{1}{3}z^{2}.$
The odd-parity term in $\mathrm{d}^{2}(\theta,z)$ with $\cos\theta$ does not contribute to the integral over $\int_{0}^{\pi}\sin\theta\ d\theta$. Averaging over $z$ must include weighting by the volume of a spherical shell of radius $z$ and thickness $dz$ within a sphere of radius unity and volume $4\pi/3$:
$\widebar{\mathrm{d}^{2}}=3\int_{0}^{1}z^{2}\left[1-\frac{1}{3}z^{2}\right]\,dz=\frac{4}{5}.$ (20)
@@ Appendix C Estimation of the effect of forward scattering on the mean free path
When neutrons of around 1 MeV energy elastically scatter from a highAA nucleus, the scattering has a narrow forward-directed peak, with a forward-to-backward ratio in the differential cross section of order 10 to 1. Forward-directed scattering does not contribute significantly to the diffusive random-walk process. Historically, the elastic cross section was also given as a transport cross section, which includes an additional factor $\langle 1-\cos\theta\rangle$. Nowadays it is rare to see the transport cross section listed with neutron scattering parameters, however, it can be estimated as follows.
@@ Appendix D Application to the diffusion equation
$D\left[\nabla_{r}n(r)\right]\Big|_{r=R}=\frac{1}{4}n_{g}\widebar{v}$ (29)
@@
$\frac{\dot{N}}{N}=\frac{1}{\tau_{w}}=\frac{\widebar{v}}{\ell}=\beta\frac{1}{(4/3)\pi R^{3}}\ \frac{\widebar{v}}{4}\ 4\pi R^{2}$ (33)
@@
$\ell^{\prime}+\frac{4}{3}x+\frac{6}{\pi}x^{2}=0$ (36)